# I’M HOI: Inertia-aware Monocular Capture of 3D Human-Object Interactions

### v1

@@ Abstract
We are living in a world surrounded by diverse and “smart” devices with rich modalities of sensing ability. Conveniently capturing the interactions between us humans and these objects remains far-reaching. In this paper, we present I’m-HOI, a monocular scheme to faithfully capture the 3D motions of both the human and object in a novel setting: using a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). It combines general motion inference and category-aware refinement. For the former, we introduce a holistic human-object tracking method to fuse the IMU signals and the RGB stream and progressively recover the human motions and subsequently the companion object motions. For the latter, we tailor a category-aware motion diffusion model, which is conditioned on both the raw IMU observations and the results from the previous stage under over-parameterization representation. It significantly refines the initial results and generates vivid body, hand, and object motions. Moreover, we contribute a large dataset with ground truth human and object motions, dense RGB inputs, and rich object-mounted IMU measurements. Extensive experiments demonstrate the effectiveness of I’m-HOI under a hybrid capture setting. Our dataset and code will be released to the community at afterjourney00.github.io/IM-HOI.
{strip}
Figure 1: Taking a monocular RGB video and single inertial measurement unit (IMU) sensor recording, our approach, I’m-HOI achieves efficient and robust capture results on challenging and dynamic human-object interaction (HOI) motions like skateboarding.
@@ 1 Introduction
In I’m-HOI, our key idea is adopt a two-stage paradigm to make full use of both the object-mounted IMU signals and the RGB stream, which consists of general motion inference and category-aware motion refinement. For the former stage, we introduce holistic human-object tracking in an end-to-end manner. Specifically, we generate human motions via a multi-scale CNN-based network for 3D keypoints, followed by an Inverse Kinematics (IK) optimization layer. To reason the companion object motions, we progressively fuse the human features with IMU measurements via object-orientated mesh alignment feedback. We also adopt a robust optimization to refine the tracked object pose and improve the overlay performance, especially when the object is invisible in the RGB input. For the second refinement stage, we propose to tailor the conditional motion diffusion models [21, 43] for utilizing category-level interaction priors. During training the diffusion model corresponding to a certain object, we treat the tracked motions and the raw IMU measurements from the previous stage as the condition information. We also adopt a novel over-parameterization representation with extra regularization designs to jointly consider the body, object, and especially hand regions during the denoising process. Thus, our refinement stage not only projects the initial human-object motions onto the category-specific motion manifold but also infills possible hand motions for vividly capturing human-object interactions. To train and evaluate our I’m-HOI, we contribute a large multi-modal dataset of human-object interactions, covering 295 interaction sequences with 10 diverse objects in total 892k frames of recording. We also provide ground truth body, hand, and 3D object meshes, with dense RGB inputs and rich object-mounted IMU measurements. To summarize, our main contributions include:
@@
  • We adopt an efficient holistic human-object tracking to progressively fuse the motion features, companion with a conditional diffusion model to refine and generate vivid interaction motions.
@@
Figure 2: The pipeline of I’m-HOI. Assuming video and inertial measurements input, our approach consists of a general interaction motion inference module (Sec. 3.1) and an category-specific interaction diffusion filter (Sec. 3.2) to capture challenging interaction motions.
@@ 2 Related Work
Human motion capture from single RGB images and videos has achieved tremendous progress [5, 50, 89, 31, 53, 56, 75, 39, 38, 36, 40, 66, 37, 100, 42, 103, 101] since the parametric body model SMPL [47, 68, 57] was released. However, contextual human-object and human-scene interactions (HOIs/HSIs) reconstruction from monocular input is far more challenging. The pioneer work PHOSA [104] proposes a purely optimization-based framework to estimate static human-object spatial arrangements relying on handcrafted contact heuristics, which is unscalable and error-prone to depth ambiguities. Benefited from emerging 3D interaction motions datasets [69, 15, 109, 7, 16, 4, 24, 22, 83, 12, 26], learning-and-optimization work [85, 25] give promising results by modeling human-object relative distance field in data-driven manner and joint post-optimization. The state-of-the-art video-based method VisTracker [86] further incorporates motion infilling techniques [96] to enable space-time coherent tracking. However, these approaches still suffer from unacceptable runtime cost and unsatisfying accuracy under complex interaction scenarios.
@@
Complementary to vision-based methods, human motion capture using inertial measurement units (IMUs) has also been extensively studied. Previous commercial solutions [87, 52] can capture accurate and detailed motion with dense sensors. Since the exploration of SIP [80], data-driven methods under sparse sensors configuration [23, 93, 94, 28, 84] realize real-time performance, deriving consumer-level products [72]. To mitigate the inherent defects of single-modal systems, multi-modal approaches [18] fuse inertial signals with RGB [61, 62, 79, 48, 78, 81, 13, 20, 110, 49, 30, 46, 54], RGBD [19, 112], ego-view [95], and LiDAR [67] references, achieving balanced local pose estimation and global localization. In this work, we extend multi-sensor fusion strategy to 3D human-object interactions capture, which is validated profitable for both accuracy and efficiency.
@@
Human motion priors have been demonstrated crucial for realism of capture and synthesis on multiple modeling methodologies, including predefined kinematic structure [2, 113], GMM [57], GAN [14, 31, 3], VAE [34, 66, 105, 59], MLP [77] and more cutting-edge, diffusion models [21, 71, 76, 70, 97, 32]. Additionally, context-aware human motion synthesis [106, 111, 51] and scene placement generation [91, 92] successfully extract contextual prior knowledge from data. More recently,  [58, 88, 44] model dynamic interaction patterns but ignore object category-level distribution differences. On another track,  [108, 41] focus on interactions with chairs, which is static and lacks diversity. This work aims to learn object category-specific interaction prior to model dynamic interaction distributions between human and diverse objects.
@@ 3 Method
We present a new paradigm for 3D human-object interactions capture in a lightweight and hybrid setting: using a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). As illustrated in Figure 2, we propose a general interaction motion inference module (Sec. 3.1) to jointly recover human-object spatial arrangements in end-to-end manner. An category-specific interaction diffusion filter (Sec. 3.2) is tailored to refine capture results from the former with the learned object category-level prior.
@@ 3.1 General Interaction Motion Inference
Existing vision-based methods [85, 86] follow fitting-learning-optimization framework, which, in our observation, is fragile to heavy or prolonged human-object occlusions, inefficient in inference time cost and poor in generalization ability as reported in Sec. 5.2. In contrast, we take object as an extra body joint, and propose to estimate human-object spatial arrangement in a holistic and end-to-end manner. An optional optimization procedure can be implemented to further improve capture accuracy.
@@
Given a monocular image sequence $\bm{I}\in\mathbb{R}^{T\times h\times w\times 3}$, we first segment human and object mask $\bm{S}_{h},\bm{S}_{o}\in\mathbb{R}^{T\times h\times w\times 1}$ separately using SAM [35]. Following that, a pre-trained ResNet-34 [17] image encoder is adopted to extract image feature from stacked RGB and object mask. After that, We take the raw inertial rotation $\bm{Q}\in\mathbb{R}^{T\times 6}$, acceleration $\bm{A}\in\mathbb{R}^{T\times 3}$ and normalized object template $\mathcal{O}$, combined with $\bm{I},\bm{S}_{o}$ as our network input. Our approach outputs human shape $\bm{\beta}\in\mathbb{R}^{T\times 10}$ and pose $\bm{\theta}\in\mathbb{R}^{T\times 3N_{J_{b}}}$, object rotation $\bm{R}_{o}\in\mathbb{R}^{T\times 6}$ and translation $\bm{T}_{o}\in\mathbb{R}^{T\times 3}$. Here, $T$ is the sequence length, $h\times w$ is the resolution of images and masks, $N_{J_{b}}=22$ is the number of body joints. We adopt standard SMPL model [47] for human motion representation.
@@
We first develop a multi-scale CNN-based network to detect 3D human body joints $\bm{J}\in\mathbb{R}^{T\times 3N_{J_{b}}}$ and object center together. Given the extracted image feature, we feed it into a group of deconvolution layers appended with a final convolution layer to reconstruct 3D keypoint heatmaps. Then we take the expectation of each heatmap to determine 3D keypoint positions [74]. Note that all the keypoints are canonicalized to root-relative representation. Commonly used combination of 3D keypoints and 2D reprojection loss $\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}$ is taken to train the CNN network, and simultaneously fine-tune the image extractor. Following 3D keypoint estimation, an off-the-shelf pre-trained inverse kinematics layer proposed in [103] is adopted and fine-tuned by $\mathcal{L}_{\text{twist}}$ to recover human motion $\bm{\theta}$ from estimated $\hat{\bm{J}}$. Please refer to [85, 86, 42, 103] for detailed loss functions.
For object tracking, we can get an initially-posed object through $\mathcal{C}(\bm{Q})\mathcal{O}+\hat{\bm{T}_{o}}$ with estimated object translation and raw rotation data from the object-mounted IMU sensor, where $\mathcal{C}(\cdot)$ is the mapping from 6D rotation representation [114] to rotation matrix. In order to eliminate systematic biases in $\bm{Q}$ and correct inaccurate prediction $\hat{\bm{T}_{o}}$ under occlusions, we attach one MLP-based regressor to each of the intermediate multi-scale image feature grid [100, 101], which forms a feedback loop to estimate corrective increment of object motion progressively. In the $i$-th loop, we first uniformly sample $N_{S}=400$ vertices on the posed object mesh $\mathcal{C}(\hat{\bm{R}}_{o}^{(i)})\mathcal{O}+\hat{\bm{T}}_{o}^{(i)}$. Then, we project vertices onto the $i$-th feature map to obtain object mesh-aligned feature, which is subsequently fed into the $i$-th regressor to predict $\Delta\hat{\bm{R}}_{o}^{(i)}$ and $\Delta\hat{\bm{T}}_{o}^{(i)}$. Particularly, $\hat{\bm{R}}_{o}^{(0)}=\mathcal{C}(\bm{Q})$. The training loss of the feedback network group is defined as $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$, where $\mathcal{L}_{\text{occ-sil}}$ is the occlusion-aware silhouette loss proposed in [104]. We find that better results can be achieved with an augmented silhouette area loss:
$\mathcal{L}_{\text{area}}=\sum_{i=0}^{N_{F}-1}||\sum\mathcal{D}(\hat{\bm{R}}_{o}^{(i)}\mathcal{O}+\hat{\bm{T}}_{o}^{(i)})-\sum\bm{S}_{o}||_{2}^{2},$ (1)
where $\mathcal{D}(\cdot)$ refers to differentiable rendering function [65] and $N_{F}=3$ is the number of feedback iterations. Starting from here, we re-define $\hat{\bm{R}}_{o}=\hat{\bm{R}}_{o}^{N_{F}-1}$ and $\hat{\bm{T}}_{o}=\hat{\bm{T}}_{o}^{N_{F}-1}$.
@@
To improve object tracking accuracy, especially in invisible cases, we propose an optional optimization module. In addition to visual cues, we further constraint object rotation and trajectory to inertial measurements. The energy function is formulated as:
@@
Specifically, $\mathcal{E}_{\text{visual}}$ minimizes the discrepancy between the rendering result and the segmentation of object: $\mathcal{E}_{\text{visual}}=\sum||\mathcal{D}(\hat{\bm{R}}_{o}\mathcal{O}+\hat{\bm{T}}_{o})-\bm{S}_{o}||_{2}^{2}$. At the same time, $\mathcal{E}_{\text{imu}}$ regularizes object motion temporally:
@@ 3.2 Category-specific Interaction Diffusion Filter
For the second refinement stage, a category-specific interaction motion filter is proposed to (i) project capture results from the previous stage onto the manifold; (ii) infill hand motions conditioned on body-object interaction motions.
@@
[…]
@@
In reverse process, we formulate condition information as a tuple $\bm{c}=(\bm{j}_{h_{b}},\bm{j}_{o},\bm{\theta}_{h_{b}},\bm{\theta}_{o},\bm{q},\bm{a})\in\mathbb{R}^{216}$ and concatenate it with masked hand motion $\bm{m}=(\bm{j}_{h_{h}},\bm{\theta}_{h_{h}})\in\mathbb{R}^{270}$, where $\bm{j}_{h_{b}}\in\mathbb{R}^{66}$, $\bm{\theta}_{h_{b}}\in\mathbb{R}^{132}$ means body-only joint positions and rotations. We follow [64] to predict $\bm{x}_{0}$ itself as $\hat{\bm{x}}_{\phi}(\bm{x}_{n},n,\bm{c})$, where $\phi$ represents the parameters of neural network. The training loss is $L1$-norm simple objective [21, 43]:
@@
Specifically, $\mathcal{L}_{\text{off}}$ enforces the predicted object center that lies in a small region determined by the offsets relative to 52 body-hand joints as:
$\mathcal{L}_{\text{off}}=\sum_{i=0}^{N_{J}}||(\hat{\bm{j}}_{o}-\hat{\bm{j}}_{h}^{(i)})-(\bm{j}_{o}-\bm{j}_{h}^{(i)})||_{1}.$ (8)
We then constraint the reproduced human body-hand joints consistent with the body model skinned by predicted joint rotations:
@@
where $\mathcal{M}(\cdot)$ refers to forward function of SMPL-H and $\mathcal{J}(\cdot)$ is the joint regressor. In order to temporally smooth human motion, the velocity term $\mathcal{L}_{\text{vel}}$ is formulated as:
@@
where $\bm{j}_{t}=[\bm{j}_{h,t},\bm{j}_{o,t}]$. Finally, $\mathcal{L}_{\text{imu}}$ guides the generated object poses and the trajectory conform with IMU measurements, which improves robustness under invisible scenarios:
$\mathcal{L}_{\text{imu}}=\mathcal{L}_{\text{rot}}+\mathcal{L}_{\text{acc}}.$ (11)
@@
$\mathcal{L}_{\text{rot}}=||\hat{\bm{\theta}}_{o}-\bm{q}||_{1},$ (12)
@@
$\mathcal{L}_{\text{acc}}=||(\hat{\bm{j}}_{o,t}-\hat{\bm{j}}_{o,t-1}+\frac{\bm{a}_{t}\tau^{2}}{2})-(\bm{j}_{o,t+1}-\bm{j}_{o,t})||_{1},$ (13)
where $\tau$ is the time interval between two consecutive frames. It is noted that, related work [86] simulates such second-order constraints in a pseudo manner [99] to eliminate mutation of first-order signals. In contrast, we incorporate acceleration explicitly [46, 67].
[…]
@@
Table 1: Quantitative comparison with several baselines on both human and object tracking accuracy.

### v2

@@ Abstract
We are living in a world surrounded by diverse and “smart” devices with rich modalities of sensing ability. Conveniently capturing the interactions between us humans and these objects remains far-reaching. In this paper, we present I’m-HOI, a monocular scheme to faithfully capture the 3D motions of both the human and object in a novel setting: using a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). It combines general motion inference and category-aware refinement. For the former, we introduce a holistic human-object tracking method to fuse the IMU signals and the RGB stream and progressively recover the human motions and subsequently the companion object motions. For the latter, we tailor a category-aware motion diffusion model, which is conditioned on both the raw IMU observations and the results from the previous stage under over-parameterization representation. It significantly refines the initial results and generates vivid body, hand, and object motions. Moreover, we contribute a large dataset with ground truth human and object motions, dense RGB inputs, and rich object-mounted IMU measurements. Extensive experiments demonstrate the effectiveness of I’m-HOI under a hybrid capture setting. Our dataset and code will be released to the community at the project page.
{strip}[Uncaptioned image]
Figure 1: Taking a monocular RGB video and a single inertial measurement unit (IMU) sensor recording, our approach, I’m-HOI, efficiently and robustly captures challenging and dynamic human-object interactions (HOI), such as skateboarding.
Corresponding author
@@ 1 Introduction
In I’m-HOI, our key idea is to adopt a two-stage paradigm to make full use of both the object-mounted IMU signals and the RGB stream, which consists of general motion inference and category-aware motion refinement. For the former stage, we introduce holistic human-object tracking in an end-to-end manner. Specifically, we generate human motions via a multi-scale CNN-based network for 3D keypoints, followed by an Inverse Kinematics (IK) optimization layer. To reason the companion object motions, we progressively fuse the human features with IMU measurements via object-orientated mesh alignment feedback. We also adopt a robust optimization to refine the tracked object pose and improve the overlay performance, especially when the object is invisible in the RGB input. For the second refinement stage, we propose to tailor the conditional motion diffusion models [21, 44] for utilizing category-level interaction priors. During training the diffusion model corresponding to a certain object, we treat the tracked motions and the raw IMU measurements from the previous stage as the condition information. We also adopt a novel over-parameterization representation with extra regularization designs to jointly consider the body, object, and especially hand regions during the denoising process. Thus, our refinement stage not only projects the initial human-object motions onto the category-specific motion manifold but also infills possible hand motions for vividly capturing human-object interactions. To train and evaluate our I’m-HOI, we contribute a large multi-modal dataset of human-object interactions, covering 295 interaction sequences with 10 diverse objects in total 892k frames of recording. We also provide ground truth body, hand, and 3D object meshes, with dense RGB inputs and rich object-mounted IMU measurements. To summarize, our main contributions include:
@@
  • We adopt an efficient holistic human-object tracking method to progressively fuse the motion features, companion with a conditional diffusion model to refine and generate vivid interaction motions.
@@
Figure 2: The pipeline of I’m-HOI. Assuming video and inertial measurements input, our approach consists of a general interaction motion inference module (Sec. 3.1) and a category-specific interaction diffusion filter (Sec. 3.2) to capture challenging interaction motions.
@@ 2 Related Work
Since the release of the parametric body model SMPL [48, 68, 58], there has been tremendous progress [5, 51, 89, 32, 54, 57, 75, 40, 39, 37, 41, 66, 38, 100, 43, 103, 101] in human motion capture from single RGB images and videos. However, reconstructing contextual human-object and human-scene interactions (HOI/HSI) from monocular input is far more challenging. The pioneer work PHOSA [104] proposes a purely optimization-based framework to estimate static human-object spatial arrangements relying on handcrafted contact heuristics. This approach is unscalable and error-prone to depth ambiguities. Benefited from emerging 3D interaction motion datasets [69, 15, 109, 7, 16, 4, 24, 22, 84, 12, 26], learning-and-optimization work [85, 25] has shown promising results by modeling human-object relative distance field in data-driven manner, followed by joint post-optimization. The state-of-the-art video-based method, VisTracker [86], further incorporates motion infilling techniques [96] to enable space-time coherent tracking. However, these approaches still suffer from unacceptable runtime costs and unsatisfying accuracy under complex interaction scenarios.
@@
Complementary to vision-based methods, human motion capture using inertial measurement units (IMUs) has also been extensively studied. Previous commercial solutions [87, 53] can capture accurate and detailed motion with dense sensors. Since the exploration of SIP [81], data-driven methods under sparse sensors configurations [23, 93, 94, 28, 79] have been developed to achieve real-time performance, deriving consumer-level products [72]. To address the limitations of single-modal systems, multi-modal approaches [18] fuse inertial signals with RGB [62, 63, 80, 49, 78, 82, 13, 20, 110, 50, 31, 47, 55], RGBD [19, 112], ego-view [95], and LiDAR [67] references, achieving balanced local pose estimation and global localization. In this work, we extend the multi-sensor fusion strategy to 3D human-object interactions capture, which is beneficial for both accuracy and efficiency.
@@
Human motion priors have been demonstrated crucial for realism in the capture and synthesis by multiple modeling methodologies, including predefined kinematic structure [2, 113], GMM [58], GAN [14, 32, 3], VAE [35, 66, 105, 60], MLP [77] and more cutting-edge, diffusion models [21, 71, 76, 70, 97, 33]. Additionally, context-aware human motion synthesis [106, 111, 52] and scene placement generation [91, 92] successfully extract contextual prior knowledge from data. More recent work [59, 88, 45] modeled dynamic interaction patterns but ignored object category-level distribution differences. Other methods [108, 42] focus on specific interactions with chairs, which are static and lack diversity. This work aims to learn object category-specific interaction prior to model dynamic interaction distributions between human and diverse objects.
@@ 3 Method
We present a new paradigm for 3D human-object interactions capture in a lightweight and hybrid setting: utilizing a minimal amount of RGB camera and object-mounted Inertial Measurement Unit (IMU). As illustrated in Figure 2, we propose a general interaction motion inference module (Sec. 3.1) to jointly recover human-object spatial arrangements in an end-to-end fashion. An category-specific interaction diffusion filter (Sec. 3.2) is tailored to refine capture results from the former with the learned object category-level prior.
@@ 3.1 General Interaction Motion Inference
Current vision-based methods [85, 86] typically adhere to fitting-learning-optimization framework, which we have observed to be susceptible to substantial or prolonged human-object occlusions, inefficient in inference time and limited in generalization capabilities, as discussed in Sec. 5.2. In contrast, we treat the object as an additional body joint and propose to estimate human-object spatial arrangements holistically and end-to-end. An optional optimization procedure can be incorporated to enhance capture accuracy further.
@@
Given a monocular image sequence $\bm{I}\in\mathbb{R}^{T\times h\times w\times 3}$, we first segment human and object mask $\bm{S}_{h},\bm{S}_{o}\in\mathbb{R}^{T\times h\times w\times 1}$ separately using SAM [36]. Following that, a pre-trained ResNet-34 [17] image encoder is adopted to extract image feature from stacked RGB image and object mask. After that, We take the raw inertial rotation $\bm{Q}\in\mathbb{R}^{T\times 6}$, acceleration $\bm{A}\in\mathbb{R}^{T\times 3}$ and normalized object template $\mathcal{O}$, combined with $\bm{I},\bm{S}_{o}$ as our network input. Our approach outputs human shape $\bm{\beta}\in\mathbb{R}^{T\times 10}$ and pose $\bm{\theta}\in\mathbb{R}^{T\times 3N_{J_{b}}}$, object rotation $\bm{R}_{o}\in\mathbb{R}^{T\times 6}$ and translation $\bm{T}_{o}\in\mathbb{R}^{T\times 3}$. Here, $T=64$ is the sequence length, $h\times w$ is the resolution of images and masks, $N_{J_{b}}=22$ is the number of body joints. We adopt standard SMPL model [48] for human motion representation.
@@
We first introduce a multi-scale CNN-based network to jointly detect 3D human body joints $\bm{J}\in\mathbb{R}^{T\times 3N_{J_{b}}}$ and the object center. Leveraging the extracted image feature, we feed it into a series of deconvolution layers followed by a final convolution layer to reconstruct 3D keypoint heatmaps. The 3D keypoint positions are then determined by the expectation of each heatmap [74], with all keypoints canonicalized to root-relative representation except for the root joint. Commonly used combination of 3D keypoints and 2D reprojection loss $\mathcal{L}_{\text{kp3d}}+\lambda_{\text{j2d}}\mathcal{L}_{\text{j2d}}$ is utilized to train the CNN network, and simultaneously fine-tune the image extractor. Subsequent to 3D keypoint estimation, we employ and fine-tune an off-the-shelf pre-trained inverse kinematics layer [103] by $\mathcal{L}_{\text{twist}}$ to recover human pose $\bm{\theta}$ and shape $\bm{\beta}$ based on $\hat{\bm{J}}$. Please refer to [85, 86, 43, 103] for detailed loss functions.
For object tracking, we get an initially-posed object through $\mathcal{C}(\bm{Q})\mathcal{O}+\hat{\bm{T}_{o}}$ with estimated object translation and raw rotation data from the object-mounted IMU sensor, where $\mathcal{C}(\cdot)$ is the mapping from 6D rotation representation [114] to rotation matrix. In order to eliminate systematic biases in $\bm{Q}$ and correct inaccurate $\hat{\bm{T}_{o}}$ under occlusions, we attach one MLP-based regressor to each intermediate image feature grid [100, 101], which forms a feedback loop to estimate corrective increment of object motion progressively. In the $i$-th loop, we first uniformly sample $N_{S}=400$ vertices on the posed object mesh $\mathcal{C}(\hat{\bm{R}}_{o}^{(i)})\mathcal{O}+\hat{\bm{T}}_{o}^{(i)}$. Then, we project sampled vertices onto the $i$-th feature map to obtain object mesh-aligned feature, which is subsequently fed into the $i$-th regressor to predict $\Delta\hat{\bm{R}}_{o}^{(i)}$ and $\Delta\hat{\bm{T}}_{o}^{(i)}$. Particularly, $\hat{\bm{R}}_{o}^{(0)}=\mathcal{C}(\bm{Q})$. The training loss of the feedback network group is defined as $\mathcal{L}_{\text{maf}}=\lambda_{\text{occ-sil}}\mathcal{L}_{\text{occ-sil}}+\lambda_{\text{area}}\mathcal{L}_{\text{area}}$, where $\mathcal{L}_{\text{occ-sil}}$ is the occlusion-aware silhouette loss proposed in [104]. We find that better results can be achieved with an augmented silhouette area loss:
$\mathcal{L}_{\text{area}}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{N_{F}-1}||\sum\mathcal{D}(\hat{\bm{R}}_{o,t}^{(i)}\mathcal{O}+\hat{\bm{T}}_{o,t}^{(i)})-\sum\bm{S}_{o,t}||_{2}^{2},$ (1)
where $\mathcal{D}(\cdot)$ refers to differentiable rendering function [29] and $N_{F}=3$ is the number of feedback iterations. Here, we re-define $\hat{\bm{R}}_{o}=\hat{\bm{R}}_{o}^{N_{F}-1}$ and $\hat{\bm{T}}_{o}=\hat{\bm{T}}_{o}^{N_{F}-1}$.
@@
It’s noteworthy that human mask $\bm{S}_{h}$ is only required in $\mathcal{L}_{\text{occ-sil}}$ during training but not when testing.
@@
To improve object tracking precision, especially in invisible cases, we propose an optional optimization module. In addition to visual cues, we further constraint object rotation and trajectory to inertial measurements. The energy function is formulated as:
@@
Specifically, $\mathcal{E}_{\text{visual}}$ minimizes the discrepancy between the rendering result and the segmentation of object: $\mathcal{E}_{\text{visual}}=\frac{1}{T}\sum_{t=0}^{T-1}\sum||\mathcal{D}(\hat{\bm{R}}_{o,t}\mathcal{O}+\hat{\bm{T}}_{o,t})-\bm{S}_{o,t}||_{2}^{2}$. At the same time, $\mathcal{E}_{\text{imu}}$ regularizes object motion temporally:
@@ 3.2 Category-specific Interaction Diffusion Filter
In the second refinement stage, a category-specific interaction motion filter is proposed to (i) project capture results from the preceding stage onto the manifold; (ii) infill hand motions conditioned on body-object interaction motions.
@@
Figure 3: We exhibit selected highlights of IMHD2 on the left side, and 10 well-scanned objects on the right side. In total, our dataset comprises 295 sequences and captures approximately 892k frames of data.
@@
In reverse process, we formulate condition information as a tuple $\bm{c}=(\bm{j}_{h_{b}},\bm{j}_{o},\bm{\theta}_{h_{b}},\bm{\theta}_{o},\bm{q},\bm{a})\in\mathbb{R}^{216}$ and concatenate it with masked hand motion $\bm{m}=(\bm{j}_{h_{h}},\bm{\theta}_{h_{h}})\in\mathbb{R}^{270}$, where $\bm{j}_{h_{b}}\in\mathbb{R}^{66}$, $\bm{\theta}_{h_{b}}\in\mathbb{R}^{132}$ represents body-only joint positions and rotations. We follow [65] to predict $\bm{x}_{0}$ itself as $\hat{\bm{x}}_{\phi}(\bm{x}_{n},n,\bm{c})$, where $\phi$ represents the parameters of neural network. The training loss is $L1$-norm simple objective [21, 44]:
@@
Specifically, $\mathcal{L}_{\text{off}}$ enforces the predicted object center to lie in a small region determined by the distance offsets relative to 52 body-hand joints as:
$\mathcal{L}_{\text{off}}=\frac{1}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{N_{J}}||(\hat{\bm{j}}_{o,t}-\hat{\bm{j}}_{h,t}^{(i)})-(\bm{j}_{o,t}-\bm{j}_{h,t}^{(i)})||_{1}.$ (8)
We then constraint the reproduced human body-hand joints consistent with the body model skinned from predicted joint rotations:
@@
where $\mathcal{M}(\cdot)$ refers to forward function of SMPL-H model and $\mathcal{J}(\cdot)$ is the joint regressor. In order to temporally smooth human motion, the velocity term $\mathcal{L}_{\text{vel}}$ is formulated as:
@@
where $\bm{j}_{t}=[\bm{j}_{h,t},\bm{j}_{o,t}]$. Finally, $\mathcal{L}_{\text{imu}}$ guides the generated object poses and the trajectory conform to IMU measurements, which improves robustness under invisible scenarios:
$\mathcal{L}_{\text{imu}}=\mathcal{L}_{\text{rot}}+\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}_{\text{acc},t}.$ (11)
@@
$\mathcal{L}_{\text{rot}}=\frac{1}{T}\sum_{t=0}^{T-1}||\hat{\bm{\theta}}_{o,t}-\bm{q}_{t}||_{1},$ (12)
@@
$\mathcal{L}_{\text{acc},t}=||(\hat{\bm{j}}_{o,t}-\hat{\bm{j}}_{o,t-1}+\frac{\bm{a}_{t}\tau^{2}}{2})-(\bm{j}_{o,t+1}-\bm{j}_{o,t})||_{1},$ (13)
where $\tau$ is the time interval between two consecutive frames. It’s noteworthy that, related work [86] simulates such second-order constraints in a pseudo manner [99] to eliminate mutation of first-order signals. In contrast, we incorporate acceleration explicitly [47, 67].
Figure 4: Qualitative 3D capturing results of I’m-HOI on IMHD2 dataset. Each sample includes an RGB image input, captured motion from camera view, and side-view visualization.
IMHD2 (fast) BEHAVE (slow to medium) [4] InterCap (slow) [24]
@@
Table 1: Quantitative comparison was conducted with several baselines on both human and object tracking accuracy.